40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 58 of 349  Not logged in ELOG logo
ID Date Author Type Category Subject
  14719   Tue Jul 2 16:57:09 2019 gautamUpdateCDSc1sus is flaky

Since the work earlier this morning, the fast c1sus model has crashed ~5 times. Tried rebooting vertex FEs using the reboot script a few times, but the problem is persisting. I'm opting to do the full hard reboot of the 3 vertex FEs to resolve this problem.

Judging by Attachment #1, the processes have been stable overnight.

Attachment 1: c1sus_timing.png
c1sus_timing.png
  14718   Tue Jul 2 12:30:53 2019 gautamUpdateElectronicsAcromag crate switched to Sorensens

[chub, gautam]

We crossed off another couple of bullets today.

It took me ~1 hour to realize that c1susaux requries the running of sudo /sbin/ifup eth0 to be run in order to see the martian network - why???

Activity:

  1. Stopped the c1susaux machine:
    • Moved alignment sliders of ITMX and ITMY to 0 as a precaution.
    • Shutdown the c1susaux machine so that it doesn't become unhappy with the missing Acromags when we power the unit down.
  2. Dialled down supply voltages on the +/- 15 V and +/- 20 V DC Sorensens. Current draw became 0 A on the front panel indicators.
  3. Chub tapped some new terminal blocks for +15 V DC and +20 V DC
    • This required some additional daisy chaining, which is why we dialled down the Sorensens.
    • New cables were made using the "standard" LIGO color scheme, which isn't really applicable in this case because we are using +15 V DC (orange sheath wire) and + 20 V DC (yellow sheath wire) whereas the closest LIGO standard voltages are +18 V DC and +24 V DC.
    • A test cable, presumably meant to be used in the electronics area (orange for +15 V DC) was destroyed for this work as we opted for speed rather than making a new cable.
  4. Disconnected bench power supplies that were powering the Acromags, and connected the new cables.
    • I opted to use 5 A fuses in the terminal blocks for these supplies as the current draw is pretty significant.
  5. Dialled the Sorensens back up to the nominal voltages:
    • Attachment #1 shows the front panels of the Sorensens before and after this work.
    • The current limit on the +20 V DC Sorensen had to be raised, because the Acromag box draws ~2.3 A on its own, whereas the previous current draw was 2.8 A.
  6. Brought the c1susaux machine back online. Took me a while to get to the bottom of why I wasn't able to see c1susaux on the martian, but eventually, I figured out the whole sbin/ifup thingy. 

I don't understand the exact chain of causation, but during this work, the fast c1sus model crashed. I had to go through a few iterations of the scripted vertex machine rebooting, but things seem to be back in a normal state now, see Attachment #2. Should probably run the IFO test suite to make sure everything is a-okay, but for now, I am able to lock the IMC so I'm moving on.

The main task remaining here is to take new pictures of everything and upload to the wiki. Also, need to update the Sorensen labels to reflect their current values, some of them are outdated.

Quote:
  • Take photos of the new setup, cabling.
  • Remove the old c1susaux crate from the rack to free up space, possibly put the PSL monitoring acromag chassis there.
  • Test that the OSEM PD whitening switching is working for all 8 vertex optics.(verified as of 5/3/19 5pm)
  • New 15V and 24V power cables with standard LIGO connectors need to be run from the Sorensenn supplies in 1X5. The chassis is currently powered by bench supplies sitting on a cart behind the rack.
  • All 24 new DB-37 signal cables need to be labeled.
  • New 96-pin DIN connectors need to be put on two ribbon cables (1Y5_80 B, 1Y5_81) in the 1X4 rack. We had to break these connectors to remove them from the back of the eurcrates.
  • General cleanup of any cables, etc. left around the rack. We cleaned up most things this evening.
  • Rename the host computer c1susaux2 --> c1susaux, and update the DNS lookup tables on chiara.
Attachment 1: 1X5Sorensens.pdf
1X5Sorensens.pdf
Attachment 2: CDS_20190702.png
CDS_20190702.png
  14717   Tue Jul 2 12:30:44 2019 MilindUpdateComputer Scripts / ProgramsPMC autolocker

Just finished a raw version of the autolocker!! Tested it once and was able to achieve lock! This is a python version of the code at /opt/rtcds/caltech/c1/scripts/PSL/PMC/AutoLock.sh.

The current code lives in my users directory. Gautam asked me to put the completed autolocker at /opt/rtcds/caltech/c1/scripts/PSL/PMC/ and that I needn't necessarily put it on git. However, I had previously added it to my Non-linear control repo. Not sure if I should take it off? The current script still lacks some checks like those that enable it to stop after a certain time of attempting to lock or those that handle interrupt signals. Will do that in some time.

P.S. As Koji says, Victory! :-P

P.P.S. Rana pointed out that this is not the objective and what we actually wanna do is run a search over the parameter space of the locking process. I will document my ideas about this process as soon as I do a little more reading. He also said that it would not do to have command line arguments as the main source from which parameters are procured and that .yml files ought to be used instead. I will make that change asap.

 

Quote:

I've begun working on this. Steps to complete:

  1. Convert the autolocker to python. Test that it works.
  2. Run the script with different settings of the servo gain adjust and DC output adjust parameters and obtain a plot of the average time of lock to determine what the best settings of the aforementioned parameters are.
Quote:

As Rana asked me to in the last meeting, I dug through the elogs to determine what had become of the previous autolockers. I stumbled upon this elog by Rana from before Gautam cleaned up the medm screen. Out of curiosity, I ran the autolocker script using the instructions in Rana's elog. I did this a total of 5 times and could lock the PMC 3 times fairly quickly. I attempted to decipher the details of the code but did not make much headway owing to my unfamiliarity with the language. From what I could make out from the medm screen while the autolocker was running, it appeared to be the same method as that in this elog. I will take a look at it again tomorrow. However, I intend to spend most of tomorrow working on preprocessing the data, developing the CNN script and then the simulation. 

Quote:
 
  1.  I shall also begin working on a script to autolock the PMC based on what Rana showed me on Monday. I will also take a look at the the contents of this elog and try to pick up from there. I hope to make significant progress by the next lab meeting.
  14716   Mon Jul 1 20:27:44 2019 gautamUpdateASCASX tuning

Summary:

To practise the dither alignment servo tuning, I decided to make the ASX system work again (mainly because it has fewer DoFs and so I thought it'd be easier to manage). Setup is: dither PZT mirrors on EX table-->demodulate green transmission at the dither frequencies-->Servo the error signals to 0 by an integrator.

Details:

  1. Started by checking the dither lines are showing up with good SNR in GTRX. They are, see Attachment #1. The dither lines are at 18.23 Hz, 27.13 Hz, 53.49 Hz and 41.68 Hz, and all of them show up with SNR ~100.
  2. Hand-aligned the beam till I got a maximum of GTRX ~ 0.35. This is lower than the usual ~0.5 I am used to - possibilities are (i) in the process of plugging in the BNC cable to the rear of the EX laser for my PLL investigations, I disturbed the alignment into the SHG crystal ever so slightly and I now have less green light going into the cavity or (ii) there is an iris on the EX table just before the green beam goes into the vacuum on which it is getting clipped. IIRC, I had centered the GTRX camera view such that the spot was well centered in the field of view, but now I see substantial mis-centering in pitch. So the cavity alignment for IR could also be sub-optimal (although I saw TRX ~1.15). Anyways, I decided to push on.
  3. Introduced a deliberate offset in a given DoF, e.g. M1 PIT. Then I looked at the demodulated error signals (filtered through an RLP0.5 filter post demodulation, so the 2f component should be attenuated by 100 dB at least), and tuned the demod phase until most of the signal appeared in the I-phase, which is what is used for servoing. The Q-phase signals were ~x10 lower than their I-phase counterparts after the tuning.
  4. Checked the linearity of the error signal in response to misalignment of a given DoF. I judged it to be sufficiently linear for all four DoFs about the quadratic part of the GTRX variation.
  5. Tweaked the overall servo gains to have the error signals be driven to 0 in ~10 seconds.
  6. There was quite significant cross-coupling between the DoFs - why should this be? I can understand the PIT->YAW coupling because of imperfect mounting of the PZT mounted mirror in a rotational sense, but I don't really understand the M1->M2 coupling.
  7. Nevertheless, the servo appears to work - see Attachment #2.

The adjusted demod phases, servo gains were saved to the .snap file which gets called when we run the "DITHER ON" script. Also updated the StripTool template.

I plan to repeat similar characterization on the IR dither alignment servos. I think the tuning of the ASS settings can be done independently of figuring out the mystery of why the TRY level is so low.

Attachment 1: ASX_ditherlines.pdf
ASX_ditherlines.pdf
Attachment 2: ASX.png
ASX.png
  14715   Mon Jul 1 20:18:01 2019 MilindUpdateComputer Scripts / ProgramsPMC autolocker

I've begun working on this. Steps to complete:

  1. Convert the autolocker to python. Test that it works.
  2. Run the script with different settings of the servo gain adjust and DC output adjust parameters and obtain a plot of the average time of lock to determine what the best settings of the aforementioned parameters are.
Quote:

As Rana asked me to in the last meeting, I dug through the elogs to determine what had become of the previous autolockers. I stumbled upon this elog by Rana from before Gautam cleaned up the medm screen. Out of curiosity, I ran the autolocker script using the instructions in Rana's elog. I did this a total of 5 times and could lock the PMC 3 times fairly quickly. I attempted to decipher the details of the code but did not make much headway owing to my unfamiliarity with the language. From what I could make out from the medm screen while the autolocker was running, it appeared to be the same method as that in this elog. I will take a look at it again tomorrow. However, I intend to spend most of tomorrow working on preprocessing the data, developing the CNN script and then the simulation. 

Quote:
 
  1.  I shall also begin working on a script to autolock the PMC based on what Rana showed me on Monday. I will also take a look at the the contents of this elog and try to pick up from there. I hope to make significant progress by the next lab meeting.
  14714   Mon Jul 1 20:11:34 2019 MilindUpdateCamerasSimulation enhancements

Today, I read a lot more about BRDF and modelling but could not make much headway regarding the implementation in the simulation. I've stopped for now and I'll take a crack at it tomorrow again.

Quote:

Yesterday, Rana asked me to look at Hiro Yamamoto's docs on the DCC to improve the simulation. I'm performing a first pass (=> Just skimming through to see if they're relevant, I will go through them more carefully soon!) and putting up stuff here for future reference. @Kruthi's help much appreciated!

  14713   Mon Jul 1 11:02:05 2019 MilindUpdateIOOPMC and IMC locked again, some MEDM maintenance

Made changes as discussed in this issue. Further, I need to add signal handling capabilities to the code. I belive Jon has pointed me to some code. I will take a look at that ASAP.

Quote:

> For channels corresponding to continuous values (such as say exposure time or the like) changes to abs(1+current_value)

Why abs? Is the current_value is like -5.4321 (for example for the alignment slider), this returns +4.4321 and the suspension will suffer from huge motion (well it will be returned to the original value soon though). 

  14712   Sun Jun 30 23:52:09 2019 KojiUpdateIOOPMC and IMC locked again, some MEDM maintenance

> For channels corresponding to continuous values (such as say exposure time or the like) changes to abs(1+current_value)

Why abs? Is the current_value is like -5.4321 (for example for the alignment slider), this returns +4.4321 and the suspension will suffer from huge motion (well it will be returned to the original value soon though). 

  14711   Sun Jun 30 22:21:07 2019 MilindUpdateIOOPMC and IMC locked again, some MEDM maintenance

Wrote the script. It currently lives at /users/milind/NonlinearControl/milind/unstick/unstick.py. Also pushed it to the repo here. It does the following:

  1. When run as python unstick.py c1aux (for instance) from the terminal, it parses the autoBurt.req file at /cvs/cds/caltech/target/c1aux/autoBurt.req and obtains the channels.
  2. Iterates through the channels and changes it to "some value"
    1. For channels corresponding to buttons on MEDM screen (Enable/Disable etc.) toggles between 0 and 1
    2. For channels corresponding to continuous values (such as say exposure time or the like) changes to abs(1+current_value)
  3. Resets to original value and then moves to the next channel

I tried print statements instead of actually writing to the channels as Gautam asked me to do that with supervision. Is this good enough?

Quote:

a useful piece of code that we should ask one of this summer's SURFs to write:

  1. read in a BURT .req file which is usually used to make the autosnap / restore.
  2. change ALL of the values to some value (slightly different from its current value)
  3. restore it to its current value

this will solve the sticky slider problem and do it in a systematic way. We can run it from the command line: e.g. 'unsticky.py c1psl c1ioo c1lsc'

Quote:

Aaron complained to me earlier that the PMC could not be locked. Turned out to be a classic sticky slider problem,

  14710   Sun Jun 30 22:02:26 2019 MilindUpdateCamerasKeyed c1aux crate

I wanted to try out the unstick.py script on c1aux but kept running into timeout errors. I was also confronted by a blank GigE screen. Further, couldn't telnet into c1aux using telnet c1aux as described here. Therefore, I went in and keyed the c1aux crate (1Y1).

  14709   Sun Jun 30 19:47:09 2019 ranaUpdateIOOIMC WFS agenda

we are thinking of doing a sprucing up of the input mode cleaner WFS (sensors + electronics + feedback loops)

  1. WFS Heads
    1. it has been known since ~2002 that the RF circuits in the heads oscillate. 
    2. in the attached PDF you can see that 2 opamps (U3 & U4; MAX4106) are used to amplify the tank circuit made up of the photodiode capacitance and L6.
    3. due to poor PCB layout (the output of U4 runs close to the input of U3) the opamps oscillate if the if the Reed relay (RY2) is left open (not attenuating)
    4. we need to remove/disable the relay
    5. also remove U3 for each quadrant so that it has a fixed gain of (TBD) and a 50 Ohm output
    6. also check that all the resonances are tuned to 1f, 2f, & 3f respectively
  2. Demod boards
  3. DC quadrant readouts
  4. Whitening
  5. Noise budget of sensors, including electronics chain
  6. diagonalization of sensors / actuators
  7. Requirements -
  8. Optical Layout
  9. What does the future hold ?

  1. what is our preferred pin-for-pin replacement for the MAX4106/MAX4107? internet suggests AD9632. Anyone have any experience with it? The Rabbott uses LMH6642 in the aLIGO WFSs. It has a lower slew rate than 9632, but they both have the same distortion of ~ -60 dB for 29.5 MHz.
  2. the whole DC current readout is weird. Should have a load resistor and go into the + input of the opamp, so as to decouple it from the RF stuff. Also why such a fast part? Should have used a OP27 equivalent or LT1124.
  3. LEMO connectors for RF are bad. Wonder if we could remove them and put SMA panel mount on there.
  4. as usual, makes me feel like replacing with better heads...and downstream electronics...
Attachment 1: WFS-Head.pdf
WFS-Head.pdf
  14708   Sat Jun 29 03:11:18 2019 KruthiUpdateCamerasCCD Calibration

Finding the gain of the Photodiode: The three-position rotary switch of the photodiode being used (PDA520) wasn't working, so I determined its gain by making a comparative measurement between ophir power meter and the photodiode. The photodiode has a responsitivity of 0.34 A/W at 1064 nm (obtained from the responsitivity curve given in the spec sheet using a curve digitizing software). Using the following equation, I determined the gain setting, which turned out to be 20dB.

\large Transimpedance\ Gain (V/A) = \frac{Photodiode\ reading (V)}{Ophir\ reading (W) * Responsitivity (A/W)}

Setup: Here a 1050nm (closest we have to 1064nm) LED is used as the light source instead of a laser to eliminate the effects caused by coherence of a laser source, which might affect our radiometric calibration. The LED is placed in a box with a hole of diameter 5mm (aperture angle = 40 degrees approx.). Suitable lenses are used to focus the light onto a white paper, which is fixed at an arbitrary angle and serves as a Lambertian scatterer. To make a comparative measurement between the photodiode (PDA520) and GigE, we need to account for their different sensor areas, 8.8mm (aperture diameter) and 3.7mm x 2.8 mm respectively . This can be done by either using an iris with a common aperture so that both the photodiode and GigE receive same amount of light , or by calculating the power incident on GigE using the ratio of sensor areas and power incident on the photodiode (here we are using the fact that power scattered by Lambertian scatterer per unit solid angle is constant). 

Calibration of GigE 152 unit: I took around 50 images, starting with an exposure time of 2000 \LARGE \mu s in steps of 2000, using the exposure_variation.py code. But the code doesn't allow us to take images with an exposure time greater than 100 ms, so I took few more images at higher exposures manually. From each image I subtracted a dark image (not in the sense of usual CCD calibration, but just an image with same exposure time and no LED light). These dark images do the job of usual dark frame + bias frame and also account for stray lights. A plot of pixel sum vs exposure time is attached. From a linear fit for the unsaturated region, I obtained the slope and calculated the calibration factor.

Equations:      \LARGE Power (P)=\frac{Photodiode\ reading(V)}{Transimpedance\ gain (V/W) * Responsivity (A/W)}                    \LARGE Calibration factor (CF) = \frac{P}{slope}

Result: CF = 1.91x 10^-16 W-sec/counts  Update: I had used a wrong value for the area of photodiode. On using 61.36 mm^2 as the area, I got 2.04 x 10^-15 W-sec/counts.

I'll put the uncertainities soon. I'm also attaching the GigE spectral response curve for future reference.

Attachment 1: calibration_setup.jpg
calibration_setup.jpg
Attachment 2: CCD_calibration_2.jpeg
CCD_calibration_2.jpeg
Attachment 3: GigE_spectral_response_curve.png
GigE_spectral_response_curve.png
Attachment 4: 152_calibration_plot.png
152_calibration_plot.png
  14706   Thu Jun 27 20:48:22 2019 MilindUpdateCamerasConvolutional neural networks for beam tracking

And finally, a network is trained!

Result summary (TLDR :-P) : No memory was used. Model trained. Results were garbage. Will tune hyperparameters now. Code pushed to github.

 

More details of the experiment:

Aim:

  1. To train a network to check that training occurs and get a feel for what the learning might be like.
  2. To set up the necessary framework to perform mulitple experiments and record results in a manner facilitating comparison.
  3. To track beam spot motion.

What I did:

  1. Set up a network that learns a framewise mapping as described in here.
  2. Training data: 0.9 x 1791 frames. Validation data: 0.1 x 1791 frames. Test data (only prediction): all the 1791 frames
  3. Hyperparameters: Attachment #1
  4. Did no tuning of hyperparameters.
  5. Compiled and fit the models and saved the results.

 

What I saw

  1. Attachment #2: data fed to the network after pre-processing - median blur + crop
  2. Attachment #3: learning curves.
  3. Attachment #4: true and predicted motion. Nothing great.

What I think is going wrong-

  1. No hyperparameter tuning. This was only a first pass but is being reported as it will form the basis of all future experiments.
  2. Too little data.
  3. Maybe wrong architecture.

Well, what now?

  1. Tune hyperparmeters (try to get the network to overfit on the data and then test on that. We'll then know for sure that all we probably need is more data?)
  2. Currently the network has around 200k parameters. Maybe reduce that.
  3. Set up a network that takes as input (one example corresponding to one forward pass)  a bunch of frames and predicts a vector of position values that can be used as continuous data).
Quote:

I got to speak to Gabriele about the project today and he suggested that if I am using Rana's memory based approach, then I had better be careful to ensure that the network does not falsely learn to predict a sinusoid at all points in time and that if I use the frame wise approach I try to somehow incorporate the fact that certain magnitudes and frequencies of motion are simply not physically possible. Something that Rana and Gautam emphasized as well.

 
Quote:
 
  1. Network training for beam spot tracking: I will begin training the convolutional network with the data pre-processed as described above. I will also simultaneously prepare data acquired from the GigE and train networks on that. Note: I planned to experiment with framewize predictions and hence did some of the work described above. However, I will restrict the number of experiments on that and perform more of those that use 3D convolution. Rana also pointed out that it would be interesting to have the network output uncertainity in the predictions. I am not sure how this can be done, but I will look into it.

 

 

Attachment 1: readme.txt
Experiment file: train.py
batch_size: 32
dropout_probability: 0.8
eta: 0.0001
filter_size: 19
filter_type: median
initializer: xavier
num_epochs: 50
activation_function: relu
dense_layer_units: 64
... 10 more lines ...
Attachment 2: frame0.pdf
frame0.pdf
Attachment 3: Learning_curves.png
Learning_curves.png
Attachment 4: Motion.png
Motion.png
  14705   Thu Jun 27 14:28:12 2019 gautamUpdateLSCPOX and POY locking

After a more systematic alignment effort, I was able to get the spots better centered on the optics (judged by eye from the analog camera views). TRY ~0.7, TRX~1.15. The X-arm dither alignment system seems to work out-of-the-box with the existing settings, I was able to run it and maximize the X-arm transmission.

Other work: I also cleaned up the area around MC2 a litte - laptop from on top of the vacuum chamber was removed and a rogue ethernet cable was also removed. The resulted in some misalignment of the IMC, which I corrected by manual alignment. Now the IMC is locked again with nominal transmission levels.

On the PSL table, I re-routed the RF output from the BeatMouth to the regular IR-ALS electronics chain (it was hijacked for PLL investigations). At EX, I disconnected the cable running from the LB1005 to the EX NPRO laser PZT (again was being used for PLL locking), and re-connected the output from the Green uPDH box to allow for some ALS tests to be done. I could then lock the EX green beam to the X-arm, and achieved GTRY ~ 0.35 using the ASX system. More to follow on ALS tests later today.

  14704   Wed Jun 26 21:01:26 2019 gautamUpdateLSCPOX and POY locking

Now that the IMC is remaining locked for extended periods of time, the next problem to attack is the ASS dither alignment system. For a start, I decided to try and get the POX and POY locking working again, as we have not fully recovered the interferometer alignment after the most recent pumpdown. I spent a couple of hours tweaking the alignment of the arm cavity mirrors, BS, and TTs to try and recover the maximum possible TRX and TRY - however, my best efforts only yielded TRX~0.8, TRY~0.75. Moreover, the beam axis is such that the spot is significantly off in YAW on both ETMs, as evidenced by the camera views (also true but less obvious on the ITMs). However, trying to bring the beam back to the center of the optics yields TRY and TRX values lower than the above reported maxima. The EX green beam is currently unavailable to verify the arm cavity alignment because of my hijacking the EX NPROs PZT control for PLL investigations, but with the Y arm, I'm able to lock a TEM00 mode. Probably just needs more careful systematic alignment, but I'm not pursuing this tonight.

  14703   Wed Jun 26 20:45:03 2019 gautamUpdateCamerasField of view options

For the beam spot position tracking, I am wondering if there is any benefit to going for a wider field of view and getting the OSEMs in the frame? It may provide some "anchor points" against which whatever algorithm can calibrate the spot position against. But there are also several point scatterers visible in the current view, and perhaps the Gaussiam beam profile moving over them and tracking the scattered intensity from these point scatterers serves the same function? I don't know of a good solution to have a "switchable" field of view configuration in the already cramped camera enclosure though.

Also, I think it may be useful to have a cron job take a picture of MC2 and archive it (once a week? or daily?) to have some long term diagnostic of how the scattered light received by the camera changes over several months.

Quote:

The GigE is focused now and I have closed the lid. I'm attaching a picture of the MC2 beam spot, captured using GigE at an exposure time of 400µs

  14702   Wed Jun 26 19:12:00 2019 KruthiUpdateCamerasGigE

The GigE is focused now (judged by eye) and I have closed the lid. I'm attaching a picture of the MC2 beam spot, captured using GigE at an exposure time of 400µs.

What was the solution to resolving the flaky video streaming during the alignment process????

-> I think, the issue was with either the poor wireless network conection or the GigE-PoE ethernet cable.

Quote:

Turns out, focusing the GigE is actually a bit tricky. With pylon, everytime I change the exposure or the focus, I'm running into the error I had mentioned earlier in one of my elogs; so I tried using the python scripts to interact with the GigE. But whenever I try to change the focal plane distance by rotating the lens coupler, the ethernet cable connection becomes loose and the camera server needs to be relaunched every now and then. Also, everytime we want to change the distance between the lenses, the telescope needs to be dismantled and refocused again. I'll try to come up with a better telescope design for this.

Yesterday, I had focused the GigE using a low exposure time and small aperture of iris, to make sure that we are actually seeing a sharp image of the beam spot. I'm attaching a picture of the beam spot I had clicked while focusing it, unfortunately, I forgot to take a picture after I had focused it completely. I'm also attaching a picture of the final setup for future reference. 


Yesterday night, Rana asked me to lock the MC2. I figured that the PSL shutter was closed; I just opened it and was able to see the beam spot on the analog camera screen.

Attachment 1: MC2_GigE_image.pdf
MC2_GigE_image.pdf
  14701   Wed Jun 26 18:28:24 2019 ranaUpdateIOOPMC and IMC locked again, some MEDM maintenance

a useful piece of code that we should ask one of this summer's SURFs to write:

  1. read in a BURT .req file which is usually used to make the autosnap / restore.
  2. change ALL of the values to some value (slightly different from its current value)
  3. restore it to its current value

this will solve the sticky slider problem and do it in a systematic way. We can run it from the command line: e.g. 'unsticky.py c1psl c1ioo c1lsc'

Quote:

Aaron complained to me earlier that the PMC could not be locked. Turned out to be a classic sticky slider problem,

  14700   Wed Jun 26 11:11:40 2019 MilindUpdateIOOPMC and IMC locked again, some MEDM maintenance

After helping Aaron key the crate and do a burt restore, I realized that it would probably be best to record the steps that Koji showed me to do a burt restore as a reference for (anyone) the future

Commands (in terminal):

  1. burttoday: changes to the directory with snapshots for the day (/opt/rtcds/caltech/c1/burt/autoburt/today)
  2. burtgooey: opens a new window with several buttons of which "Restore" needs to be selected. This opens up a second window as shown in Attachment #1. Click on Snapshot files and navigate to the snapshot you wish to restore (these are present at /opt/rtcds/caltech/c1/burt/autoburt/snapshots) and select that. A green "OK" button indicates if the Restore can be performed without a hitch. Hit "Restore" to perform the burtrestore.

 

Also Gautam explained today that the sticky slider problem is a hardware issue. That it basically means that the signal (voltage output for instance) that you request from the medm screen is not what the hardware delivers. Twice now, we have got around that with a burtrestore. My understanding of a burt restore is that it is a restoration of values from a certain time to the EPICS channels. Therefore, I don't understand why a restoration at the software level should fix how the hardware responds? Why does this happen?

Attachment 1: burtgooey.pdf
burtgooey.pdf
  14699   Wed Jun 26 10:55:13 2019 aaronUpdateIOOPMC and IMC locked again, some MEDM maintenance

The PMC was locking again after Gautam's steps above. However, after I added the directional coupler between the mixer I and the servo card (coupled to the Agilent analyzer), the PMC was again not locking, except occasionally with gain of -10 dB.

I removed the coupler (so the mixer I goes directly to the PMC servo card, as Gautam had it), and the PMC was still not locking. While checking connections, I noticed that one of the SMA cables between the LO and the mixer was not even finger tight, so I tightened them to approximately the right torque with a non-torque wrench.

This did not lead to the PMC locking, so Millind helped me key the c1psl VME crate. I burt restored the latest snapshot. Now, the PMC locks up until gain of -5. I try burt restoring the previous snapshot, which was from when the PMC was locking, and now it locks. Adding in the directional coupler again leads to the PMC not locking, though this time removing the coupler restores the normal behavior. I also tried using the coupler with the coupling port connected to a 50 Ohm terminator, and this configuration also did not lock.

I had been using a ZFDC-20-5-S+ (0.1-2000 MHz) with SMA ports and SMA-to-BNC on the input and output ports (since the mixer has BNC connectors). To reduce the number of potentially flaky connections, I am trying the ZFDC-20-4 (1-1000 MHz) that I found with BNC ports. The PMC still doesn't lock.

To get some spectrum, I've connected the PMC servo card's 'mixer out' to the Agilent's A channel, and collected a spectra from [10 Hz, 75 kHz], [75 kHz, 750 kHz], and [750 kHz, 2 MHz].


Wed Jun 26 15:23:37 2019

After the lab cleaning, I added a BNC T on the mixer I port, so now the configuration is:

Mixer I -> BNC T

-> PMC Servo card FP1TEST

-> directional coupler -> coupled to the spectrum analyzer, out port is terminated with 50 Ohms.

I thought maybe the issue was that the TF from in->out on the directional coupler is not what I expect (and Gautam suggested the in-out port might block DC), but the PMC still does not lock in the above configuration, in which the coupler is not between the mixer and the servo board--so only reflections from the coupler should matter, I think.

However, even when I plug the mixer directly into the servo board, the PMC is not locking (again) with gain above -8 dB or so. I did a burt restore again, and this fixed the problem. I wasn't sure why this burt restore is working, because all I am changing is the DC output adjust voltage and the gain, and switching on/off FP1TEST. However, I observed that after running the PMC autolocking script, observing that the autolocker did not achieve lock as it swept through resonance, and cancelling the autolocker, the PMC again cannot be locked for high gains. When I let the autolocker complete, this doesn't happen, so probably I'm just not letting some channel return to its nominal value after being changed by the autolocker.

Now after another burt restore, I'm avoiding using the autolocker and am still having trouble locking with the BNC T + directional coupler configuration above. However, now I'm noticing that the PZT control mon is always railed, as long as FP1TEST is in the loop (and independent of the output adjust voltage). I try returning to the 'baseline' configuration (mixer -> PMC servo card directly), and the PMC locks but with only 0.68 V transmission (was >0.7 V before).


Per Gautam's earlier suggestion, I switched to using the Agilent 41800A probe instead of the directional coupler. I was able to lock the OMC with this probe on a BNC T coming out of the mixer (transmission is 0.71 V). I recorded the spectra of the PMC servo board's "Mixer Out" channel, and the mixer's I as seen by the probe. I recorded spectra from 10 Hz to 100 MHz. The soft linked netgpibdata folder I had in my users directory is no longer soft linked--presumably intentional so I don't tamper with it?

I'm a bit skeptical that I've used the probe correctly, so I'm checking out the manual.

Indeed, I needed to pull back the sheath; I also noticed that the GPIB script I've been using doesn't save the data from both channels when I take a spectrum in dual mode, so I'm taking the spectra again one at a time (lights are on, IMC is locked).

  14698   Tue Jun 25 23:52:37 2019 MilindUpdateCamerasSimulation enhancements

Yesterday, Rana asked me to look at Hiro Yamamoto's docs on the DCC to improve the simulation. I'm performing a first pass (=> Just skimming through to see if they're relevant, I will go through them more carefully soon!) and putting up stuff here for future reference. @Kruthi's help much appreciated!

  14697   Tue Jun 25 22:14:10 2019 MilindUpdateCamerasConvolutional neural networks for beam tracking

I discussed this with Gautam and he asked me to come up with a list of signals that I would need for my use and then design the data acquisition task at a high level before proceeding. I'm working on that right now. We came up with a very elementary sketch of what the script will do-

  1. Check the MC is locked.
  2. Choose an exposure value.
  3. Choose a frequency and amplitude value for the applied sinusoidal dither (check warning by Gabriele below).
  4. Apply sinusoidal dither to optic.
  5. Timestamping: Record gpstime, instantaneous channel values and a frame. These frames can later be put together in a sequence and a network can be trained on this. (NEED TO COME UP WITH SOMETHING CLEVERER THAN THIS!)

Tomorrow I will try and prepare a dummy script for this before the meeting at noon. Gautam asked me to familiarize myself with the awg, cdsutils (I have already used ezca before) to write the script. This will also help me do the following two tasks-

  1. IFO test scripts that Rana asked me to work on a while ago
  2. The PMC autolocker scripts that Rana asked me work on
Quote:
 

Upcoming work (in the order of priority):

  1. Data acquisition: With the mode cleaner being locked and Kruthi having focused on to the beam spot, I will obtain data for training both GANs and the convolutional networks. I really hope that some of the work done above can be extended to the new data. Rana suggested that I automate this by writing a script which I will do after a discussion with Gautam tomorrow.

 

I got to speak to Gabriele about the project today and he suggested that if I am using Rana's memory based approach, then I had better be careful to ensure that the network does not falsely learn to predict a sinusoid at all points in time and that if I use the frame wise approach I try to somehow incorporate the fact that certain magnitudes and frequencies of motion are simply not physically possible. Something that Rana and Gautam emphasized as well.

I am pushing the code that I wrote for

  1. Kruthi's exposure variation - ccd calibration experiment
  2. modified camera_client_movie.py code (currently at /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon)
  3. interact.py (to interact with the GigE in viewing or recording mode) (currently at /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon)

to the GigEcamera repository.

 

Gautam also asked me to look at Jigyasa's report and elog 13443 to come up with the specs of a machine that would accomodate a dedicated camera server.

 

Quote:
 
  1. Network training for beam spot tracking: I will begin training the convolutional network with the data pre-processed as described above. I will also simultaneously prepare data acquired from the GigE and train networks on that. Note: I planned to experiment with framewize predictions and hence did some of the work described above. However, I will restrict the number of experiments on that and perform more of those that use 3D convolution. Rana also pointed out that it would be interesting to have the network output uncertainity in the predictions. I am not sure how this can be done, but I will look into it.
  2. Cleaning up/ formalizing code: Rana pointed out that any code that messes with channel values must return them to the original settings once the script is finished running. I have overlooked this and will add code to do this to all the files I have created thus far. Further, while most of my code is well documented and frequently pushed to Github, I will make sure to push any code that I might have missed to github.
  3. Talk to Jon!: Gautam suggested that I speak to Jon about the machine requirements for setting up a dedicated machine for running the camera server and about connecting the GigE to a monitor now that we have a feed. Koji also suggested that I talk to him about somehow figuring out the hardware to ensure that the GigE clock is the same as the rest of the system.

 

 

  14696   Tue Jun 25 15:32:16 2019 gautamUpdateIOOPMC and IMC locked again, some MEDM maintenance

Aaron complained to me earlier that the PMC could not be locked. Turned out to be a classic sticky slider problem, I keyed the c1psl VME crate, and did the usual burtrestore trick. After that, I could immediately lock the PMC and IMC with the nominal gain settings.

I also looked at the wiring at the rack. An SLP250 was installed at the mixer IF output, in parallel with a 50 ohm terminator to ground. I removed this, because as Aaron pointed out, the PMC servo card "FP1 TEST" input is already 50 ohm, and has two cascaded LC filter stages immediately after to filter out the 2f component, so the extra low-pass filtering is superfluous (in any case, 250 MHz is much too high a cutoff to be using for cutting out the 2f component which will be at ~70 MHz).

Finally, in the last ~2 weeks, we have been running with the PMC servo gain of +17 (as opposed to +18 from before). The old gain is too high, as noted by Milind. But the MEDM field for this gain goes RED unless the gain is +18. I adjusted the value of the  C1:PSL-STAT_PMC_NOM_GAIN channel to +17, so that this is no longer the case. I also edited the PMC MEDM screen to get rid of my comment that the "SLOW ADC IS DEAD" for the PMC TRANS field, since I have now hooked up the PMC trans photodiode to our temporary Acromag box.

Attachment 1: PMCctrl.png
PMCctrl.png
  14695   Tue Jun 25 11:54:47 2019 KruthiUpdateCamerasGigE

Turns out, focusing the GigE is actually a bit tricky. With pylon, everytime I change the exposure or the focus, I'm running into the error I had mentioned earlier in one of my elogs; so I tried using the python scripts to interact with the GigE. But whenever I try to change the focal plane distance by rotating the lens coupler, the ethernet cable connection becomes loose and the camera server needs to be relaunched every now and then. Also, everytime we want to change the distance between the lenses, the telescope needs to be dismantled and refocused again. I'll try to come up with a better telescope design for this.

Yesterday, I had focused the GigE using a low exposure time and small aperture of iris, to make sure that we are actually seeing a sharp image of the beam spot. I'm attaching a picture of the beam spot I had clicked while focusing it, unfortunately, I forgot to take a picture after I had focused it completely. I'm also attaching a picture of the final setup for future reference. 


Yesterday night, Rana asked me to lock the MC2. I figured that the PSL shutter was closed; I just opened it and was able to see the beam spot on the analog camera screen.

 

Attachment 1: MC2_GigE.pdf
MC2_GigE.pdf
Attachment 2: Cameras_final_setup.JPG
Cameras_final_setup.JPG
  14694   Tue Jun 25 00:25:47 2019 MilindUpdateCamerasConvolutional neural networks for beam tracking

In the previous meeting, Koji pointed out (once again) that I should determine if the displacement values and frames are synchronized before training a network. Pooja did the following last time. Koji also suggested that I first predict the motion (a series of x and y coordintates) and then slide resulting plots around until I get the best match for the original motion. This is however not possible with a neural network based approach as the network learns exactly what you show it and therefore it will learn any mismatch between the labels and the frames and predict exactly that. Therefore I came up with what Koji described as "hacky" method to achieve the same using the opencv work described previously in this elog (the only addition being the application of a mask to block out the OSEMs and work only with the beam spot) .

Hacky technique to sync frames and labels:

  1. I ran the OpenCV algorithm on the data to obtain a plot for predicted motion depicted in Attachment #2. As is evident, the predicted motion is only an approximate of the actual motion and also displays a shift . However, a plot of the fourier transform of the signal (see Attachment #1) shows that the components present are the same. However, the predominant frequency component is 0.22 Hz rather than 0.2 Hz as stated by Pooja in her elog. I wonder if this is of any consequence. Therefore, this predicted motion can be slid around until it overlaps with the applied sinusoidal dither signal "well".
  2. Defining "well": I computed an error signal as the differnece between the predicted signal and the actual motion with each signal being normalized by subtracting the mean and then dividing the resulting signal by the maximum value (see Attachment #3). The lower the power of the resulting signal, the better the synchronization of the predicted and actual signal. Note: To achieve this overlap of signals, datapoints are removed from either the start or end of the signals and this effectively reduces the number of data points available for training by 36 pionts (see Attachment #4, positive and negative shifts merely indicate if the predicted signal is being moved right or left).
  3. Attachments #5,#6 show the resutls of shifting the data by 36 samples. it is evident that there is far greater overlap of the prediction and the actual values.
  4. Well, what now? I will use the mapping between labels and frames obtained by the above steps to train a neural network.

[Koji, Milind - 21/06/2019]

  1. Well, the above is fine, but why is contour detection really necessary? Why not take a weighted sum of all the pixel values (in a rectangular region obtained, say, after blocking out the OSEMs) to see what the centroid motion is? Black areas (0 pixel intensity values) will not contribute to this sum anyway. Perhaps that can be used for the sliding instead of the above (fallible!) approach, specially for cases in which the beam "spot"  is just a collection of random speckles?
    1. Something like this was done by Pooja where she computed the sum of pixel intensities in a rectangular region containing the beam spot. However, she did this for very noisy data and observed intensity variation at a frequency double that of the applied signal.
    2. Results of applying a median filter and doing the same are presented in Attachment #7. Clearly, they can't be used for this sliding task.
    3. Results of computing the weighted sum of all the coordinates (with pixel intensities as the weights are presented in Attachment #8. Clearly, for this data and for this task, the contour approach seems to be a better method. Further, these resutls just serve to prove Rana's point that such simple, unsophisticated, naive approaches will not produce desired results and therefore, shall be presented in this very context in the report that is due.
  2. The contour detection technique does not work if the beam spot is just a cokllection of speckles. In that case Koji suggested that we use a bounding convex hull instead of a contour. Alternately, for a bunch of speckles I can perform dilation to reduce it to the same problem.
  3. Using gpstime for time stamping: To determine the absolute time which a frame is grabbed. However, the time between the time being recorded and grabbing of frame needs to be determined for this which should be doable using linux/python commands.
Quote:

Worked further on this. I skimmed through a few resources to look for details of what pre-processing can be done. Here (am planning to convert all these resources, particularly those I come across for GANs into either a README on the repo or a Wiki soon) are some of the useful things I found during today's reading. The work I skimmed through today mostly pointed to the use of a median filter for pre-processing, if any is to be done. I am presently using the Sequential() API in Keras to set up the neural network. I will train it tomorrow.

 


Upcoming work (in the order of priority):

  1. Data acquisition: With the mode cleaner being locked and Kruthi having focused on to the beam spot, I will obtain data for training both GANs and the convolutional networks. I really hope that some of the work done above can be extended to the new data. Rana suggested that I automate this by writing a script which I will do after a discussion with Gautam tomorrow.
  2. Network training for beam spot tracking: I will begin training the convolutional network with the data pre-processed as described above. I will also simultaneously prepare data acquired from the GigE and train networks on that. Note: I planned to experiment with framewize predictions and hence did some of the work described above. However, I will restrict the number of experiments on that and perform more of those that use 3D convolution. Rana also pointed out that it would be interesting to have the network output uncertainity in the predictions. I am not sure how this can be done, but I will look into it.
  3. Simulation:
    1. Putting the physics in: Previously, I worked on adding point scatterers. I shall add the effect of surface roughness and incorporate the BRDF next. Just as Gautam did, Rana also reccommended that I go through Hiro Yamamoto's work to improve my understanding of this.
    2. GANs: I will put together a readme (which I will turn into a wiki later) for all the material that I am using to develop my ideas about GAN training. Currently, my understanding of GANs is that they take as input noise vectors which are fed to the generative networks which then produce the fakes. This clearly isn't the only way to do it as GANs are used for several applications such as image generation from text. I am referring to these papers to set up the necessary architecture.
  4. PMC autolocker: I will convert the existing autolocker script to python. Rana also suggested that it would be interesting to see what the best settings of the hyperparameters would be to lock the PMC the fastest. I will write a script to do that and plot a 3D surface plot of the average time taken to lock the PMC as a function of the PZT scan speed and the Servo gain to determine the optimal setting of these "hyperparameters".
  5. Cleaning up/ formalizing code: Rana pointed out that any code that messes with channel values must return them to the original settings once the script is finished running. I have overlooked this and will add code to do this to all the files I have created thus far. Further, while most of my code is well documented and frequently pushed to Github, I will make sure to push any code that I might have missed to github.
  6. Talk to Jon!: Gautam suggested that I speak to Jon about the machine requirements for setting up a dedicated machine for running the camera server and about connecting the GigE to a monitor now that we have a feed. Koji also suggested that I talk to him about somehow figuring out the hardware to ensure that the GigE clock is the same as the rest of the system.

 

Attachment 1: Spectra.pdf
Spectra.pdf
Attachment 2: normalised_comparison_y.pdf
normalised_comparison_y.pdf
Attachment 3: residue_normalised_y.pdf
residue_normalised_y.pdf
Attachment 4: error_power_sliding.pdf
error_power_sliding.pdf
Attachment 5: normalised_comparison_y.pdf
normalised_comparison_y.pdf
Attachment 6: residue_normalised_y.pdf
residue_normalised_y.pdf
Attachment 7: intensum.pdf
intensum.pdf
Attachment 8: centroid.pdf
centroid.pdf
  14693   Mon Jun 24 15:49:05 2019 aaronUpdateIOOIMC diagnostics

aI went to repeat these measurements using the mixer out channel from the servo box, and with a slower sweep for the PDH calibration.

I had trouble getting the PDH signal, here are some notes:

  • I added a 50 Ohm terminator to BNC T on the mixer box. This had been terminated before I started, but I noticed no terminator today.
  • Noticed some distortion of my driving triangle wave if I measured it on ch3 and 4 of the tektronix scope, not present on ch 1/2
  • Initially wasn't finding a signal because I was opening the loop by turning off the Test 1 switch, but this meant the mixer mon on the servo box also did not receive the PDH signal. Instead, I cut the loop with the "BLANK" switch on the PMC screen, which instead blanks out the op amp between the mixer mon and the PMC drive conditioning (so the external drive still reaches the PZT).

attachment 1 is the configuration of the PMC screen when I was trying to get some PDH signal; I did move the DC output adjust to 0V, but found that this led to the output being railed; this makes sense, the op amp at U9 has a negative bias at GND.

Rana came by and gave me some tips.

  • I'd been using the wrong servo board diagram, it should be in D1400221
  • We removed the LP filter from the mixer output (before going to FP1TEST on the servo board), since the board itself already is filtering the IF.
  • We might have observed the thermal locking? See for yourself, the trans and refl signals while sweeping the PZT drive at 5 Hz and 30 Hz respectively are in attachments 2 and 3.
  • Rather than using an SR560, I should use an RF coupler between the mixer and FP1TEST to measure the error signal spectrum. I found a ZFDC-20-5-S+ (0.1-2000 MHz) and sent an SMA cable from the coupled port to channel R of the Agilent 4395.

We finally got the PDH signal again, and I recorded the PDH signal while driving with the following settings on the Siglent function generator.

  • 1.1 Hz triangle wave, 6Vpp, -7Vdc offset, high impedance mode

I tried getting a spectrum using the coupler, the mixer mon is seeing a DC offset though and causing the PZT to rail. Will try to understand why, but in the meantime removing the coupler (still no LP filter) lets us lock the PMC again.

RXA: Kruthi thinks all of our subsequent IMC locking problems are Aaron's fault (she was quick to give him up as soon as the thumb screws were tightened...)

Attachment 1: sweep_config_updated.png
sweep_config_updated.png
  14692   Mon Jun 24 13:48:36 2019 KruthiConfigurationCDSGiada wireless connection

[Gautam, Kruthi]

This afternoon, Gautam helped me setup Giada to access the GigE installed for MC2. Unlike Paola, which was being used earlier, Giada has a better battery life and doesn't shutdown when the charger is unplugged. Gautam configured Giada to enable its wireless connection to Martian, just like Koji had configured Paola (https://nodus.ligo.caltech.edu:8081/40m/14672). We also rerouted  the ethernet cable we were using with the PoE adaptor from Netgear Switch in 1x2 to 1x6.

  14691   Mon Jun 24 11:48:35 2019 gautamUpdateIOOIMC in-loop error spectra and OLTF

Attachment #1 - In loop error spectra, measured as Koji posted end of last week.

  • Main difference is that the line noise seems much lower.
  • For the "dark" measurement, I set the IN1 gain of the servo board to the value of +4 dB, which is what it is in lock.
  • As Koji mentioned, this isn't an apple-to-apple comparison as the IMC loop will squish the plotted orange trace.
  • Nevertheless, the fact that the blue trace is above orange everywhere gives confidence that we are in fact measuring frequency noise.
  • For the higher frequency measurement, I used the AG4395 analyzer, which has 50 ohm input impedance. So to get the measurements with the SR785 to line up, I multiplied these by x2.
  • For the frequency axis calibration, I used the value of 13 kHz/V for the PDH discriminant, which was what I measured it to be last year (but I didn't check again today).
  • Note that the IMC locking loop OLTF has not been undone, so this isn't the actual laser frequency noise on the transmitted beam. In order to measure the latter, we'd have to use (for example) an arm cavity as an analyzer.

Attachment #2 - OLTF of the IMC loop.

  • Measurement was made using the IN1/IN2 method, injection was done at the "A EXC" front panel BNC input.
  • For comparison, I've overlaid a measurement from the 2017 IMC loop investigations. Doesn't seem to be significantly different.
  • UGF and phase margin are in the ballpark of what they were reported to be in the past.

Attachment #3 - Photo courtesy Koji showing the bank of BNC connectors used for these measurements.

Clearly, these measurements were taken in a time when the IMC was "well behaved". How to characterize what's happening when this isn't the case?

Attachment 1: IMCfreqNoise.pdf
IMCfreqNoise.pdf
Attachment 2: IMC_OLTF.pdf
IMC_OLTF.pdf
Attachment 3: IMC_CMboard.jpg
IMC_CMboard.jpg
  14690   Mon Jun 24 08:12:10 2019 gautamUpdateIOOIMC is locking normally again

Over the last 24 hours, the IMC autolocker was able to keep the MC locked ~60% of the time. This is not particularly good, but is an improvement on ~2 weeks ago when the IMC couldn't be locked.

There are two periods, which I've indicated by vertical cursors, between which the autolocker was doing something strange - usually this kind of trend is caused by one or more of the VME crates being unresponsive and the autolocker gets stuck, but I confirmed that both c1psl and c1iool0 are telnet-able. So I conclude that the stability and reliability of the IMC loop is still not as good as it used to be.

Note also that while the PC drive RMS level mostly hovers around 1 V, there are several excursions above that level. This in itself isn't a new phenomenon. I will do some more characterization by measuring the in-loop error signal spectrum and maybe the OLTF of the IMC locking loop.

Quote:
 

Let' see how stable or otherwise the config is. I must've jiggled some poor cable connection back into a good spot while working on the PSL table? Or the NPRO decided to be less noisy on Sunday.

Attachment 1: IMCdutycycle.png
IMCdutycycle.png
  14689   Sun Jun 23 14:43:14 2019 KojiUpdateIOOIMC is locking normally again

Note that I have removed an SR785, an oscilloscope, some SRS instruments from the PSL and PMC last night.

But they (and RF Network Analyzer) were not there when the problem started.

We should record the IMC error (at test point monitor) too? If the IMC locks on Monday too, I'll do it.

  14688   Sun Jun 23 09:36:32 2019 gautamUpdateIOOIMC is locking normally again

After typing up the elog, I decided to try locking the IMC again - now it locks again with the "OLD" gain settings. I tested it ~5 times, the autolocker brings the lock back and the PC drive levels are normal. IMC transmission and MC REFL DC light levels in lock are normal. The PC Drive RMS voltage is <1V. What's more, there is no longer any evidence of 60 Hz line harmonics any more in the PMC diagnostics channels. Compare attachment #1 to this elog.

WTF.

I undid the changes Koji made to the autolocker gains, and am trying the old settings again. Let' see how stable or otherwise the config is. I must've jiggled some poor cable connection back into a good spot while working on the PSL table?

Anyway, this helps Kruthi and Milind.

Attachment 1: PMCdiag.pdf
PMCdiag.pdf
  14687   Sun Jun 23 08:09:53 2019 gautamUpdateIOONPRO diagnostics

Summary:

Over the last few days, I've been doing some (complementary) measurements to what Aaron and Koji have been looking at. The motivation was to identify if the problems we are seeing are optical (i.e. imprinted on the PSL light) or electronic. My findings:

  1. 60 Hz line noise in PMC REFL and PMC TRANS is heavily dependent on whether I connect cables between the measuring PDs and Acromag ADC or not - but even with the Acromag cable disconnected, the 60 Hz RIN is HUGE - 10 mVpp out of 670 mV DC, and lines are much dirtier if you have connections to the SLOW ADCs. Measurement was made by looking at the time-domain signals on a battery powered Tektronix oscilloscope. See Attachment #1. I believe this line noise is higher it was. Cause is unknown to me at this point.
  2. The NPRO noise eater seems to function as advertised. The measured RIN with the noise eater enabled (our nominal operating condition) is in line with what the manual tells us it should be. See Attachment #2.
  3. There isn't strong evidence of excess frequency noise (measured with PLL) out to 100 kHz. I didn't measure the high-frequency part yet, but maybe I'm doing something wrong with the PLL setup which should be first corrected. See Attachments #3, #4.
  4. The beat note frequency between the free-running PSL and EX NPRO's is definitely slewing more than the quadrature sum of the advertised 1 MHz/min slewing per the manual.

Evidence:

Attachment #1: Time domain look at PMC Refl and Trans signals under various operating conditions. During this work, I took the chance to remove ~4 BNC T connectors that were connected on the PMC TRANS photodiode (Thorlabs). Now, there is one cable going to the Acromag ADC, and one going to the Oscilloscope used to monitor these signals. Any further T-ing can be done at the oscilloscope.

Attachment #2: RIN measurement of the NPRO light. I opted to place a Thorlabs PDA55 in the IR ALS pickoff light path. This is before the light sees the PMC. A DC block was inserted between the PDA55 and the AG4395 used to make the measurement. DC level of the PD output was 3.1 V into high-Z and I used half this value to normalize the measurement made by the 50-ohm input AG4395 into RIN units. The measurement was made with the PZT and slow temperature controls to the NPRO connected/disconnected, but I saw no significant difference. 

Attachment #3: Frequency noise measurement via PLL. This shows the loop transfer funtion for the PLL. Some details of the setup:

  • The beat note for locking the PLL was made between the PSL NPRO and the EX NPRO (output of the IR ALS BeatMouth). ~4dBm beatnote.
  • Local oscillator was sourced by a Marconi, f_carrier=33 MHz, RF level = +10dBm.
  • Level 7 Mixer and LB1005 controller from the mode-spectroscopy PLL setup.
  • PLL control signal routed to EX NPRO PZT via Heliax cable running along south arm. 
  • Why EX and not PSL or Marconi FM? Latter has limited range, ~1/10th of that offered by NPRO PZT. PSL PZT has a 2.9 Hz corner freq Pomona box. I could disconnect this for the purpose of PLL locking, but I thought it may be interesting to see if there’s any hints of the problem being electrical, by looking at PLL spectra with / without Pomona box. The expected delay due to cabling is only 400 ns, so not really a limiting factor for the PLL bandwidth.
  • LB 1005 settings:
    • PI corner = 3 kHz.
    • G = 2.30 (I could not increase this further - with the PSL+Lightwave NPRO PLL, we could achieve a UGF of ~60 kHz, but in this setup, I can't do much better than ~7kHz before the loop starts oscillating, not sure if the fact that the PZT actuation coefficient for the Innolight is ~5x lower than for the Lightwave is enough to explain this?).
    • LFGL = 90 dB.
  • Mixer output had a maximum value of 800 mVpp => PLL discriminant is 400 mV/rad.
  • The "eye fit" is just the transfer function of two poles at DC (one for frequency to phase conversion in the PLL and one for the LB1005 integrator), and a zero at 3kHz (PI corner). I scaled the gain till the "fit" and measurement lined up, and then used this model to undo the loop suppression of the error signal to extract the frequency noise without worrying about the frequency vector of the measurement being limited.
  • Once again, slow temperature control and PZT controls to the PSL NPRO were disconnected so this measurement was made with two free-running NPROs.

Attachment #4: Frequency noise measurement via PLL. This shows the frequency noise. I've overlaid the expected frequency noise between 2 free-running NPROs, model used is in the text box in the plot. There isn't strong evidence of excess high frequency noise in this measurement. The fact that the "LB 1005 input terminated" trace is below all the others supports the hypothesis that I'm measuring real frequency noise. The bump around a few kHz could indicate some gain peaking?

However, I'm unable to find good agreement between the measured frequency noise using the error point and the control point. For the former, I used the PLL discriminant mentioned above of 400 mV/rad, and undid the loop suppression, and for the latter I used a PZT discriminant of 1.7 MHz/V. However, there is still a constant scale difference between these two traces. So I'm doing something wrong?

Next steps:

  1. More interpretation of the PLL measurement results required.
  2. Measure the PLL error signal spectrum to higher frequencies using the AG4395. 
  3. ???

I've not disturbed the PLL setup in case anyone else wants to repeat these measurements, but I have restored the normal electrical connections to the PSL PZT and temperature control.

Some other activity:

  1. Alignment into the PMC was tweaked.
  2. NPRO laser pump current was increased from 1.9 A to 2.0 A.
  3. PMC servo gain was changed from +18 to +17 to prevent the servo from oscillating.
Attachment 1: consolidatedOscopeScreenCaps.pdf
consolidatedOscopeScreenCaps.pdf
Attachment 2: RINcomp.pdf
RINcomp.pdf
Attachment 3: PLL_OLTF.pdf
PLL_OLTF.pdf
Attachment 4: PLLnoise.pdf
PLLnoise.pdf
  14686   Fri Jun 21 19:36:26 2019 KojiUpdateIOOIMC diagnostics

The IMC REFL error signal was measured to compare it with the other spectra (if we have).

The blue curve is the in-loop IMC error and the red is the dark noise. So they are not an apple-to-apple comparison. But the red noise is going to be suppressed by the loop, and still the red is below blue. This means that the blue curve is the measured noise rather than the readout noise.

We suspect that the current issue is the PC drive saturation (as usual). Does this indicate that the laser freq noise is actually increased?

----

Another suspect was that the degradation of the LO level. We used to have the issue of slowly dying ERA-5 (ERA-5SM indeed). The RF levels on the demod board were measured using an active probe.

The LO input: 0dBm, ERA-5 input: -2.7dBm and -2.1dBm for I and Q. I found that the outputs of the ERA-5SM were +10.5dBm and +10.6dBm.
This lead me to replace the chips but the situation was not changed. Then I realized that the LO levels should have been measured with the load replaced from the mixers to a 50Ohm load. Somehow these mixers lower the apparent LO levels. So I decided to say this is OK.

Attachment 1: IMC_error.pdf
IMC_error.pdf
  14685   Fri Jun 21 19:22:40 2019 KojiConfigurationBHDReviving the single OMC BHD design?

I think a Faraday rotator rotates the polarizations in a same way for both forward and backward beam, and it's not like in this figure.
And the transmission through multiple faradays will also be a big issue.

  14683   Wed Jun 19 19:12:51 2019 aaronUpdateIOOIMC diagnostics

Here are the results from the fit. Data can be found on nodus in /users/aaron/40m/data/PMC/190617/. I've put a jupyter notebook with the analysis in /users/aaron/40m/analysis/PDH_calibrate.ipynb (might be some filename issues due to different directory structure on my laptop).

Here's a summary of the current measurement. I'll be referencing the diagram for the PMC servo card.

  1. With the PMC servo loop open, sweep the PMC PZT by sending a triangle wave in to J5 (external drive on the servo card).
    1. I used a 100Hz drive, but should use something slower so my drive isn't filtered out by the 100Hz pole on the servo card and the 10Hz pole on the PZT.
  2. Monitor the voltage at the HV drive, as well as at "mixer out" (J8 on the servo board)
    1. Note that I took this PDH error signal from FP2TEST rather than "mixer out", which means my error signal was not low pass filtered.
  3. Calibrate the HV mon in units of Hz by fitting the PDH signal. The sidebands should be 35.5Mhz away from the carrier peak.
    1. This part needs to be done differently to account for thermal locking in the PMC.
  4. You now have the PDH error signal as a function of PMC resonance in Hz, and can use that to calibrate the PDH error spectrum.
    1. The spectrum is taken when the PMC is locked, so the Hz/V scaling is the slope of the PDH error signal.

In the figures below, I obesrved that for fast (100Hz) drives, the PDH error signal had a pi/2 phase shift relative to the triangle wave, which means even though the resonance appears near the turnaround of the triangle, it is actually occuring near the center of the range.

There are several problems with this data:

  • PMC error signal spectrum is not properly calibrated, even according to the process described above
  • The drive was faster than the response of the PZT.
  • I was driving with a ~1V excitation, so I've lost a factor of 10 somewhere on the way to the external drive curve. Probably just a problem with how I've read the data dump from the scope.
Attachment 1: PMC_Error_Spectrum.pdf
PMC_Error_Spectrum.pdf
Attachment 2: PDH_signal.pdf
PDH_signal.pdf
Attachment 3: PDH_signal_full.pdf
PDH_signal_full.pdf
  14682   Tue Jun 18 22:54:59 2019 MilindUpdateCamerasConvolutional neural networks for beam tracking

Worked further on this. I skimmed through a few resources to look for details of what pre-processing can be done. Here (am planning to convert all these resources, particularly those I come across for GANs into either a README on the repo or a Wiki soon) are some of the useful things I found during today's reading. The work I skimmed through today mostly pointed to the use of a median filter for pre-processing, if any is to be done. I am presently using the Sequential() API in Keras to set up the neural network. I will train it tomorrow.

Quote:

Begun setting up an environment (as mentioned before, on my local machine) and scripts to run experiments with Convolutional networks for beam tracking. All code has been pushed to this folder in the GigEcamera repository. I am presently looking for pre-processing techniques for the video which go beyond the usual "Crop the images! Normalize pixel values! Convert to Grayscale!".

 

  14681   Tue Jun 18 20:35:07 2019 aaronUpdateIOOIMC diagnostics

I made a script (/users/aaron/40m/GPIB/tds_dump.py) that grabs data from a Tektronix scope and packages into a pickled dict with the following structure:

  1. ch1
    1. times ("ts")
    2. values ("vals")
    3. channel info ("info")
  2. ch2
    1. ""
  3. etc

I made a python notebook that does the following:

  1. Grab the data from the pickle above
  2. Fit a triangle wave to the drive signal
  3. Determine the (change in Volts) / second from the triangle wave, as well as define the times of a single sweep of the PDH error signal
  4. Trim the error signal data to contain the PDH signal from the carrier and two sidebands only (the original trace was for three periods).
  5. Fit the functional form of the PDH signal to the trimmed error signal. 
    1. The sideband frequency is fixed at 35.5 MHz, and the scaling of Volts-to-Hz is left free, so this fit gives the calibration of IF volts to Hz.
  6. Grab the spectra (already saved from the Agilent with the netgpib scripts) and apply this V-to-Hz scaling
  7. Plot the spectra

The fit in step (5) is still looking quite bad, despite the fitted values being close to the expected. Since we really just want a calibrated spectrum, I'll instead fit a line to the linear portion of the PDH error signal for the carrier and both sidebands, then determine the scaling from that.

  14680   Mon Jun 17 22:19:04 2019 MilindUpdateComputer Scripts / ProgramsPMC autolocker

As Rana asked me to in the last meeting, I dug through the elogs to determine what had become of the previous autolockers. I stumbled upon this elog by Rana from before Gautam cleaned up the medm screen. Out of curiosity, I ran the autolocker script using the instructions in Rana's elog. I did this a total of 5 times and could lock the PMC 3 times fairly quickly. I attempted to decipher the details of the code but did not make much headway owing to my unfamiliarity with the language. From what I could make out from the medm screen while the autolocker was running, it appeared to be the same method as that in this elog. I will take a look at it again tomorrow. However, I intend to spend most of tomorrow working on preprocessing the data, developing the CNN script and then the simulation. 

Quote:
 
  1.  I shall also begin working on a script to autolock the PMC based on what Rana showed me on Monday. I will also take a look at the the contents of this elog and try to pick up from there. I hope to make significant progress by the next lab meeting.

  14679   Mon Jun 17 16:02:17 2019 aaronUpdateComputerskeyed PSL crate

Milind pointed out that all boxes on the medm screens were white. I didn't have diagnostics from the medm screens, so I started following the troubleshooting steps on the restart procedures page.

It seemed like maybe a frontend problem. I tried telnet-ing into several of the fe, and wasn't able to access c1psl. The section on c1psl mentions that if this machine crashes, the screens will go white and the crate needs to be turned off and on. Millind did this.

Now, most of the status lights are restored (screenshot).

 


Milind: I did a burtrestore following this and locked the PMC following the steps described in this elog.

Attachment 1: after_keying_crate.png
after_keying_crate.png
  14678   Mon Jun 17 14:36:13 2019 MilindUpdateCamerasConvolutional neural networks for beam tracking

Begun setting up an environment (as mentioned before, on my local machine) and scripts to run experiments with Convolutional networks for beam tracking. All code has been pushed to this folder in the GigEcamera repository. I am presently looking for pre-processing techniques for the video which go beyond the usual "Crop the images! Normalize pixel values! Convert to Grayscale!".

Quote:

Networks for beam tracking:

  1. I will use the architectures suggested in this work with a few modifications. I will use MSE loss function, Adam optimizer and my local GPU for training.

 

  14677   Mon Jun 17 12:37:16 2019 aaronUpdateIOOIMC diagnostics

Grabbed the PMC data

I went to set up the spectrum analyzer measurements through GPIB, but inadvertently deleted the contents of ~/Agilent/netgpibdata/ (made a soft link in my folder, decided I wanted it gone but rm'ed instead of unlink). I copied what I think was in that folder back (from /opt/rtcds/caltech/c1/scripts/general/labutils/netgpibdata).

Again, the spectra are:

  • R-- PMC TRANS PD into SR560 with G=2 DC coupled, no filtering
  • A-- PDH IF into SR560 with G=2, DC coupled, no filtering
  • B-- PMC PZT HV MON into SR560 with G=2, AC coupled, no filtering

I recorded the three spectra with the following parameters:

  • Three separate spans:
    • 10 Hz to 150 kHz
    • 100 to 550 kHz
    • 500 kHz to 2.5 MHz
  • bwSpanRatio = 0.1 %
  • averages: 10
  • number of points: 801
  • spec type: noise (PSD units)

I then ran AGmeasure with the above parameters in the yaml, with the rest following the defaults in AgilentTemplate.yaml. I saved the data in /users/aaron/40m/data/PMC/190617/

Looks like the header contains all of the parameters, so I shouldn't have trouble distinguishing the spectra. I didn't get the instant plotting working, but the data seem to be there.

I'm still having trouble getting the data from the oscilloscope. I'm not sure why the tektronix scrips I've used before aren't working, I'm checking it now.

update: Grabbed the data, the issue was just using the wrong IP address. 

 

  14676   Sat Jun 15 00:03:26 2019 KruthiUpdateCamerasGigE setup

The analog camera is aligned and we are able to see all the 4 OSEMs (pictures attached). Due to secondary reflection from the beamspiltter (BS1-1064-33-2037-45S), when the MC2 is locked, we are getting a ghost image of the beam spot along with the primary image. 


The pylon app in Paola was reporting an error saying "0xE1000014: The buffer was incompletely grabbed". I followed the instructions given in this site, and changed the 'Packet Size' to 1500 and 'Inter-Packet Delay parameter' to a value greater than 20,000 (µs). This did the trick and I was able to use the continuous shot mode without any interruption. I'm attaching a picture of MC2 that I captured using GigE.

Quote:

Today, I tried aligning it further; I'm attaching a picture of it. We are not able to see all the 4 OSEMs yet. In the reference picture I had taken, before taking off the previous analog setup, the OSEMs are not seen. So, I don't really understand what the other 2 spots seen on the current screen are. Are they actually OSEMs?

I need a laptop next to MC2, so that I can have a look at it and make further alignments. So, I tried accessing the GigE attached to the telescope using Paola. The pylon app in it, throws an error, few seconds after running it in continuous shot mode, and disconnects the GigE; everything works fine on Rossa though. I'll put up further details soon.

Quote:

don't need to lock - make sure the 4 OSEMs are centered on the camera field just as we have for the arm cavity mirrors

Quote:

I'm attaching a picture of the screen. I just positioned the enclosure by turning it a bit and I suppose we can see the mirror inside the vacuum now (the MC2 is still not locked). 

 

 

 

Attachment 1: mc2_GigE.pdf
mc2_GigE.pdf
Attachment 2: MC2_analog.jpeg
MC2_analog.jpeg
Attachment 3: MC2_analog_OSEMs.jpeg
MC2_analog_OSEMs.jpeg
  14675   Fri Jun 14 13:10:00 2019 aaronUpdateIOOIMC diagnostics

The circuit diagram for the PMC servo card is D980352. From this diagram, I see that I can send an excitation from the network analyzer to FP2TEST (9.09 kOhm input impedance) where it is added to the PMC error signal before going to the loop filters.

I hook up the following

  • Agilent 4395A output to SR560 (300 Hz HP, gain of 1)
  • SR560 to FP2TEST and to Agilent's channel R
  • PMC error signal IF (mixer box mounted to rack, I noticed the IF BNC->SMA were a bit loose/stressed by a hanging LP RF filter) to SR560 (also 300 Hz HP, gain of 2)
  • SR560 to Agilent channel A

I 'Enable' Test 2 on the PSL screen, so FP2TEST gets added to the error signal.

PDH signal

  • TDS 3034B with four channels
    • 1. PMC servo box external drive (split off from the function generator)
    • 2. PMC servo box output monitor (mirrors the drive, shows when drive is saturating)
    • 3. IF signal (split off after the mixer)
    • 4. PMC Trans (long BNC from the PSL table)
  • SRS DS345 function generator (into the PMC servo box' external drive)
    • 100 Hz signal (there's a 10 Hz pole on the PZT drive, so any faster than this and I can't see both sidebands without the HV output mon clipping)
    • 3.19 Vpp amplitude (smallest amplitude at 100 Hz such that both sidebands are well resolved)
    • 1.52 V offset (center the carrier's PDH error signal at pi/2 out of phase with the drive)

I was able to see the carrier and both sidebands.

I tried to grab this data from the scope via ethernet, but was unsuccessful (timeout errors, I'm using the scripts from scripts/tektronix/tek-dump, and the GPIB box that Kruthi had been using for the GigE cam; I also tried plugging in directly scope->ethernet. Never got anything but timeout errors, so maybe I'm not specifying the port correctly. Anyway the trace is frozen on the scope for later use, or I can easily repeat this now that I know how).

Spectrum

Next, I locked the PMC (Test1 is off, tune DC output adjust until I get some transmission, turn on the loop at Test1, increase the gain to before the loop goes unstable). I'm sending the following channels to SR560 (gain = 2, no filtering, high dynamic reserve, 50 Ohm outputs), and reading spectra from the Agilent 4395A:

  • R-- PMC TRANS PD
  • A-- PDH IF
  • B-- PMC PZT HV MON

The HV mon was always saturating the preamp, so I disconnected it; I added a 50 Hz (6db) high pass to the Trans PD signal, since it has a DC component.

I got to take a look at the traces on the spectrum analyzer front panel, but too tired to do the GPIB for now. There are peaks, things look reasonable.

  14674   Fri Jun 14 00:40:33 2019 KruthiUpdateCamerasGigE setup

Today, I tried aligning it further; I'm attaching a picture of it. We are not able to see all the 4 OSEMs yet. In the reference picture I had taken, before taking off the previous analog setup, the OSEMs are not seen. So, I don't really understand what the other 2 spots seen on the current screen are. Are they actually OSEMs?

I need a laptop next to MC2, so that I can have a look at it and make further alignments. So, I tried accessing the GigE attached to the telescope using Paola. The pylon app in it, throws an error, few seconds after running it in continuous shot mode, and disconnects the GigE; everything works fine on Rossa though. I'll put up further details soon.

Quote:

don't need to lock - make sure the 4 OSEMs are centered on the camera field just as we have for the arm cavity mirrors

Quote:

I'm attaching a picture of the screen. I just positioned the enclosure by turning it a bit and I suppose we can see the mirror inside the vacuum now (the MC2 is still not locked). 

 

 

Attachment 1: Reference_image_taken_with_previous_analog_camera_setup.jpeg
Reference_image_taken_with_previous_analog_camera_setup.jpeg
Attachment 2: MC2_image.JPG
MC2_image.JPG
  14673   Thu Jun 13 22:46:41 2019 KojiUpdateIOOLeft IMC at the intermediate gains

SURFS want some locking of IMC for camera adjustment.

So I left the IMC with intermediate gains so that it keeps locking and unlocking.

VCO (overall) iMC gain of -32, FSS common gain 3, and the FAST gain 20. I believe MC2tickle is ON too.

  14672   Thu Jun 13 22:21:44 2019 KojiConfigurationCDSPaola wireless connected to martian

SURFs had trouble connecting paola to martian via wireless.
Of course, it requires a fixed IP but it had not it yet. So I went to chiara and gave 192.168.113.110 as "paolawl". Note that the wired connection has .111 and it is "paola".

Followed the instruction on http://nodus.ligo.caltech.edu:8080/40m/14121

  14671   Thu Jun 13 21:29:52 2019 MilindUpdateCamerasSteps to interact with GigE

As directed by Gautam, I have set up one script- interact.py (at /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon/interact.py) to perform the following two tasks:

  1. View the GigE feed for a fixed period of time.
  2. Record the GigE feed for a fixed amount of time.

 

Steps to view GigE feed for a fixed amount of time:

  1. Run the following commands in the terminal to navigate to the concerned directory and then view the feed
    1. cd /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon
    2. python interact.py --path_to_config <path_config> --mode 0 --view_time <viewing_time>, where
      1. path_config: full path to configuration file, defaults to /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon/C1-CAM-ETMX.ini if --path_to_config is not used
      2. viewing_time: time in seconds for which the feed is to be displayed. The server is closed  after this time and the window freezes and can be manually closed.
    3. Exiting the feed in between: The script terminates automatically after the specified time. To terminate the feed in between, close the window manually using the x icon the top right. This makes sure that the server is correctly closed. If closed using the Ctrl-C command in the terminal, the server is left running and any attempt to unwittingly set up another results in an error (see Attachment #1). In this case, the server and client processes needs to be identified manually and killed. I have used the following steps
      1. ps -eaf | grep server, then identify the PID for the python camera_server.py process
      2. kill PID
      3. similarly for the client file

Steps to record the GigE feed for a fixed amount of time:

I tried to look for elegant solutions that wouldn't require editing the code that Jon wrote and stumbled upon this useful bit of information but ended up deciding that it was just easier to change the camera_client_movie.py (/opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon/camera_client_movie.py). It can still be run as previously described, where video recording is terminated by using Ctrl-C. Steps to record for a fixed period of time are

  1. cd /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon
  2. python interact.py --path_to_config <path_config> --mode 1 --save_time <recording_time> --file_name filename, where\
    1. path_config: full path to configuration file, defaults to /opt/rtcds/caltech/c1/scripts/GigE/SnapPy_pypylon/C1-CAM-ETMX.ini if --path_to_config is not used
    2. recording_time: time in seconds for which the feed is to be saved. No video is displayed during this time.
    3. filename: full path to the file where the video is to be saved. Overwrites any existing files.

I'll make aliases for these to make the whole process more user friendly. I'm halting this for now and will discuss what else needs to be done once Gautam gets back.


Regarding the autolocker: I spoke to Aaron today and as he is in tomorrow, I'll ask him about the burt files and the ideal configuration.

I'm also starting with GANs now.

 

 

Attachment 1: terminal_error_server.pdf
terminal_error_server.pdf
  14670   Thu Jun 13 18:01:18 2019 aaronUpdateIOOIMC diagnostics

Continuing this investigation of the IMC, today I am getting familiar with the PMC and FSS. I'd like to measure the frequency noise of the PSL referenced to the PMC.

I checked that the PSL shutter is off, so no light reaches the IMC.

I'm not really sure what I'm looking for on the FSS boxes. I found a few documents to guide:

I ran the FSS autolock script from C1PSL_FSS, nothing obvious changes when I do so. The FSS error signal (which I think is PSL-FSS_MIXERM) is flatlined, and the RC-RF_PD has no LO (PSL-FSS_LOCALC is nan).

  14669   Thu Jun 13 15:08:31 2019 MilindUpdateElectronicsVCO pickup by Rich

Rich dropped by at around 3:00 PM today and picked up the VCO in Attachment #1 and left the note in Attachment #2 on Gautam's desk with the promise of bringing it back soon.

Attachment 1: WhatsApp_Image_2019-06-13_at_15.06.57.jpeg
WhatsApp_Image_2019-06-13_at_15.06.57.jpeg
Attachment 2: WhatsApp_Image_2019-06-13_at_15.06.57(1).jpeg
WhatsApp_Image_2019-06-13_at_15.06.57(1).jpeg
  14668   Thu Jun 13 14:28:46 2019 ranaUpdateCamerasGigE setup

don't need to lock - make sure the 4 OSEMs are centered on the camera field just as we have for the arm cavity mirrors

Quote:

I'm attaching a picture of the screen. I just positioned the enclosure by turning it a bit and I suppose we can see the mirror inside the vacuum now (the MC2 is still not locked). 

 

ELOG V3.1.3-